### v6

# FormulaReasoning: A Dataset for Formula-Based Numerical Reasoning

Abstract
The application of formulas (e.g., physics formulas) is a fundamental human ability in solving numerical reasoning problems. Existing numerical reasoning datasets rarely explicitly state the formulas employed, as their questions often rely on implicit commonsense mathematical knowledge. To address this gap, we introduce FormulaReasoning, a new dataset specifically designed for formula-based numerical reasoning. It consists of 5,324 questions that require numerical calculations grounded in external physics formulas. We provide normalized, fine-grained annotations in both English and Chinese, including formula structures, parameter names, symbols, numerical values, and units—curated through extensive manual effort with LLM-assisted validation to ensure high quality. Additionally, we offer a consolidated formula database to serve as an external knowledge source. We analyze various reasoning approaches on FormulaReasoning, with emphasis on comparative evaluation of different architectural and methodological frameworks. Our assessment includes retrieval-augmented methods, approaches that decompose reasoning into formula generation, parameter extraction, and numerical calculation, as well as optimization techniques using preference data. We identify key challenges in formula-based numerical reasoning that require further investigation across different reasoning paradigms, highlighting opportunities for methodological advancement.
FormulaReasoning: A Dataset for Formula-Based Numerical Reasoning
Xiao Li* and Bolin Zhuthanks: These authors contributed equally to this work. and Kaiwen Shi Sichen Liu and Yin Zhu and Yiwei Liu and Gong Chengthanks: Corresponding author State Key Laboratory for Novel Software Technology, Nanjing University, Nanjing, China {xiaoli.nju,bolinzhu}@smail.nju.edu.cn {kaiwenshi,sichenliu,yinzhu,ywliu}@smail.nju.edu.cn gcheng@nju.edu.cn
1 Introduction
Numerical reasoning constitutes one of the significant forms within natural language reasoning (Frieder et al., 2023). The study of numerical reasoning has seen substantial progress in recent years, driven largely by the development of LLMs (OpenAI, 2023; Touvron et al., 2023; Yang et al., 2024) and specialized datasets (Wang et al., 2017; Dua et al., 2019; Amini et al., 2019; Cobbe et al., 2021a). Current datasets for numerical reasoning typically include simple commonsense numerical questions that under-reflect the complexity of real-world problems. These datasets either do not provide process supervision information, or the provided reasoning steps are essentially incomplete, as they often rely on implicit commonsense knowledge but not explicit knowledge guiding the reasoning process. This issue becomes particularly evident when LLMs experience hallucination (Frieder et al., 2023; Bang et al., 2023), especially in the absence of clear knowledge guidance. Consequently, one might ask “What explicit knowledge could be used to guide a numerical reasoning process?”. Formulas exactly represent such knowledge that has been largely overlooked in previous research but is frequently utilized in real-life applications.
Figure 1: An example from FormulaReasoning. Numerical values with units given in the question and obtained from intermediate steps are highlighted in red and purple, respectively. Formulas and their elements are in blue.
Limitations of existing datasets
Take a question from the GSM8K dataset (Cobbe et al., 2021a) as an example: “A robe takes 2 bolts of blue fiber and half that much white fiber. How many bolts in total does it take?” This question implicitly uses simple commonsense mathematical knowledge (e.g., half means dividing by 2) to solve without requiring complex domain-specific formular knowledge for numerical reasoning. Recently, Liu et al. Liu et al. (2023) constructed two formula-based datasets, Math23K-F and MAWPS-F. Although 33% and 38% of the questions in these datasets, respectively, require the use of formulas, they still mainly consist of simple commonsense formulas (e.g., total_amount = unit_amount $\times$ total_number). Currently, there is a lack of datasets where questions require domain-specific formulas to guide complex numerical reasoning, such as the physics formula used to calculate the heat absorption of an object.
Our work
To fill the gap, we constructed FormulaReasoning, a new dataset for numerical reasoning that requires the use of domain-specific formulas (i.e., physics). We annotated formulas for each question as illustrated in Figure 1. This formula-based feature enables FormulaReasoning to effectively differentiate model capabilities across different scales, particularly highlighting the limitations of smaller language models in handling domain-specific formula applications. In Table 1, we compare FormulaReasoning with two existing formula-based datasets and the well-known formula-free datasets GSM8K and MATH (Hendrycks et al., 2021). Compared to Math23K-F and MAWPS-F, FormulaReasoning contains more diverse formulas (272 vs. 18–51). Moreover, the higher average number of reasoning steps (2.39 vs. 1.01–1.16) provides a more comprehensive evaluation framework for understanding how different model architectures handle multi-step formula-based reasoning.
Table 1: Statistics of Math23-F, MAWPS-F, GSM8K, MATH, and our FormulaReasoning.
Dataset Math23K-F MAWPS-F GSM8K MATH FormulaReasoning # questions 23,162 2,373 8,792 12,500 5,324 # questions requiring formulas 7,750 911 N/A N/A 5,324 # formulas (and variants) 51 (131) 18 (46) 0 (0) 0 (0) 272 (845) Avg. # reasoning steps 1.16 1.01 3.59 Not Provided 2.36
Specifically, we collected questions requiring formula-based numerical reasoning from junior high school physics exams. With combined efforts of manual annotation and LLM assistance, we annotated each question with an explanation text that provides normalized reasoning steps with relevant formulas (including formula structures, parameter names, symbols, numerical values, and units) and a final answer. We also built a consolidated formula database that functions as an external knowledge base and can be used by retrieval-based/augmented systems. The original questions and annotations are in Chinese. We used an LLM to translate them into high-quality English, forming a bilingual dataset.
We conducted extensive experiments on FormulaReasoning to compare different reasoning paradigms, including standard LLM approaches with 8B to >>100B parameters, fine-tuned models with Chain-of-Thought (CoT) supervision, data augmentation methods, and retrieval-augmented generation (RAG) methods. In addition, we derived preference data and evaluated direct preference optimization (DPO) methods. Our analysis reveals significant differences in how various approaches handle formula-based reasoning.
Our contributions are summarized as follows.
  • We construct a formula-based numerical reasoning dataset with fine-grained annotations for each question. It can be applied to evaluate knowledge-guided reasoning capabilities. Its normalized reasoning steps are useful for the emerging study of step-supervised reasoning.
  • We perform an evaluation of LLMs of various sizes, RAG, fine-tuned small models, and DPO methods based on derived preference data. Our experimental results establish a solid baseline for future research and also indicate that the performance disparity across model scales presents new research opportunities for efficient formula-based reasoning.
FormulaReasoning is available from HuggingFace https://huggingface.co/datasets/cat-overflow/FormulaReasoning. Our code is available from GitHub https://github.com/nju-websoft/FormulaReasoning.
2 Related Work
2.1 Numerical Reasoning Datasets
The study of numerical reasoning in natural language has existed for years. Numerous datasets, such as GSM8K (Cobbe et al., 2021b), TSQA (Li et al., 2021), and MATH (Hendrycks et al., 2021), have addressed natural language numerical reasoning. Another line of research that focuses on numerical reasoning in natural language is math word problem (MWP). MWP tasks typically provide a short passage (i.e., a question) and require the generation of an arithmetic expression that can compute an answer. Representative datasets include MAWPS (Koncel-Kedziorski et al., 2016), Math23K (Wang et al., 2017), MathQA (Amini et al., 2019), etc. Several works focus on numerical reasoning in specialized domains. Examples include GeoSQA (Huang et al., 2019), which focuses on the geography domain, STEM (Drori et al., 2023) and ScienceQA (Lu et al., 2022) which cover multiple disciplines in science and technology. Distinguished from these datasets, the numerical reasoning questions in our FormulaReasoning have explicitly labeled formulas.
The recently introduced datasets Math23K-F and MAWPS-F (Liu et al., 2023) require formulas for 33% and 38% of the questions, respectively, but their formulas are all simple commonsense formulas (e.g., total_cost = unit_cost $\times$ total_number). By contrast, our FormulaReasoning adapts questions from physics exams, with every question accompanied by a fine-grained annotation of domain-specific formulas. In addition, we provide a consolidated formula database that can serve as an external knowledge base to assess RAG solutions.
2.2 Numerical Reasoning Methods
Methods for numerical reasoning have evolved from statistical approaches (Hosseini et al., 2014; Kushman et al., 2014) to those based on rules and templates (Shi et al., 2015; Wang et al., 2019) and further to deep learning models (Gupta et al., 2019; Chen et al., 2022; Kim et al., 2022; Li et al., 2023a). In recent years, with the rapid development of LLMs, they have demonstrated strong capabilities to resolve numerical reasoning questions. Consequently, several methods have been proposed to improve the reasoning abilities of LLMs, including the notable CoT method (Wei et al., 2022), along with many subsequent variant approaches (Kojima et al., 2022; Wang et al., 2022; Zhou et al., 2022; Li et al., 2023b). Preference learning methods have also emerged (Rafailov et al., 2023).
Our experiments implemented and evaluated representative existing methods as baselines for FormulaReasoning, including zero-shot CoT prompting for LLMs ranging from 8B to more than 100B parameters. We trained a specialized formula retriever to be used in RAG. We divided the reasoning process into formula generation, parameter extraction, and numerical calculation, and used data augmentation to enhance fine-tuned small models. We derived preference data and performed DPO.
3 Dataset Construction
We collected questions from Chinese junior high school physics exams. We invited five graduate students to act as annotators, all of whom hold a bachelor’s degree in science and engineering. For all questions, we normalized their explanations. Each question was annotated with normalized reasoning steps in natural language and a tabular representation of the steps using formulas, including the numerical values and units for all the parameters in the formulas. This process involved a combination of manual annotation and the assistance of an LLM [^1] During dataset construction, we consistently accessed Qwen-max as our LLM via its API. to improve the efficiency of annotation. We compiled all the formulas and merged those that express the same meaning to create a consolidated formula database. Finally, we altered the questions to avoid label leakage and translated the dataset into English. In the following, we elaborate this process to construct FormulaReasoning.
3.1 Preprocessing
We crawled 18,433 junior high school physics exam questions in China from 2015 to 2024 from public sources, including only those with free-text answers and excluding multiple-choice and true/false questions. Each raw question contains a question text and an explanation text that includes reasoning steps. We eliminated questions requiring diagrams.
We filtered the questions by identifying the presence of numerical values within the explanation and confirming that the final answer was numerical. Using regular expressions, we extracted the final numerical answer including its unit from the explanation. We found that for 487 questions, our regular expressions did not return results, so we manually extracted their answers from the explanation text. After this preprocessing, we compiled an initial dataset comprising 6,306 questions.
Table 2: Original explanation and explanation with normalized formulas (highlighted in blue).
Original explanation:
[…]
Explanation with normalized formulas:
1. Calculating the temperature increase in water: [Degree of water temperature increase] = [Final temperature] - [Initial temperature] = 60 ℃ - 20 ℃ = 40 ℃. The degree of water temperature increase = 40 ℃.
2. Calculating the heat absorbed by water: [Heat absorbed by water] = [Mass of water] $\times$ [Specific heat capacity of water] $\times$ [Degree of water temperature increase] = 50 kg $\times$ 4.2 $\times$ $10^{3}$ J/(kg·℃) $\times$ 40 ℃ = 8400000 J. The heat absorbed by water = 8400000 J.
3. The thermal efficiency of the water heater can be obtained from: [Thermal efficiency of the water heater] = [Heat absorbed by water] / [Total electrical energy consumed] $\times$ 100% = 8400000 J / ($1\times 10^{7}$ J) * 100% = 84%. The thermal efficiency of the water heater = 84%.
Answer = 84%
3.2 Formula Normalization
The reasoning steps in the raw explanation text lacked a normalized format and were expressed casually. Some formulas mixed parameter names (e.g., “mass of water”) with symbols (e.g., $m_{water}$”), while others simply provided calculations over numerical values without parameter names or symbols. To ensure that all explanations adopted a common form of formulas, we performed normalization, as illustrated in Table 2. We aimed to identify and normalize the formulas in the original explanations and then verify and unify their formats. Manually carrying out such tasks would require a significant effort. However, since the process is not open-ended, but rather structured and verifiable, we could automatically, e.g. using an LLM, extract formulas from a normalized explanation, symbolicly calculate each step, and compare the result with the given answer to ensure the accuracy of our normalization.
Specifically, to improve the efficiency of annotation, we adopted a coarse-to-fine annotation process with the help of an LLM. We first prompted the LLM to revise the reasoning steps (in particular, the formulas) in the explanation text into a normalized form. Then, we prompted the LLM to correct minor errors within the normalized explanations, including formatting issues in formula annotations and inaccuracies in the parameters used during calculation. In the following, we elaborate this two-stage process. More details, including the prompts, are provided in Appendix A.1.1.
Coarse-grained annotation
We introduced each question with its original explanation and answer to guide the LLM through a few-shot prompt to normalize the explanation. We observed that the quality of the normalized explanations was generally satisfactory. We also required the LLM to present the symbol, numerical value, and unit of each parameter in formulas in the form of a table, as illustrated in Figure 1.
Fine-grained annotation
We checked the correctness of the formula format in the explanations by rules, including whether there were omissions in parameter names, symbols, or units, and these minor issues were correctable. To assess the correctness of each normalized explanation, we extracted formulas from the explanation and calculated an answer using the Numbat calculator [^2] https://numbat.dev. Numbat is designed for scientific computation with support for physical units.. If we detected that the LLM generated an ill-formatted or incorrect normalized explanation, we used a few-shot prompt to identify and correct these specific errors. We observed that the questions that remained incorrect despite multiple attempts were of inherently poor quality, e.g., missing important reasoning steps, as illustrated in Appendix A.1.2. These questions were removed from our dataset. After that, our dataset contains 5,420 questions remaining.
3.3 Formula Database Construction
Our next step was to construct a consolidated formula database for the entire dataset. Parameters in the same formula can be expressed differently in various problem contexts. For example, the two formulas “[weight of water] = [mass of water] * [gravitational acceleration]” and “[weight] = [mass] * [gravitational acceleration]” both calculate the weight of an object. It would be helpful to standardize these formulas for use with LLMs, particularly in methods such as RAG Tran et al. (2024), step supervision Zhang et al. (2024), and tree-based search Zhao et al. (2024), to avoid long-tail distribution issues caused by treating different formula expressions as different formulas.
Table 3: Changes in the number of formulas after each merging step.
Step # Formulas
Before merging 12,906
After symbolic rules based merging 1,163
After semantics based merging 439
After manual review and error correction 272
We divided the construction process of a consolidated formula database into three steps: 1) Merge formulas through symbolic rules. 2) Merge formulas through a semantics-based method. 3) Manually review and correct errors. In Table 3, we present the initial number of formulas and the remaining number of formulas after each step.
Symbolic rules based merging
This was achieved by comparing formula structures and symbols. Consider the following as an example of judging whether two formulas have the same structure. The three formulas $f_{1}:~a_{1}=(b_{1}+c_{1})/d_{1}$”, $f_{2}:a_{2}=(b_{2}+c_{2})/d_{2}$”, and $f_{3}:b_{1}=a_{1}*d_{1}-c_{1}$” have the same structure because $f_{2}$ can be derived from $f_{1}$ by renaming parameters, and $f_{3}$ can be obtained from $f_{1}$ by transformation. Moreover, in physics, physical quantities are conventionally represented by specific symbols. For example, the mass of an object is often denoted by “mm” and the density of an object is frequently represented by $\rho$”. Subscripts are then used to specify which specific object a physical quantity refers to, such as $\rho_{water}$” for the density of water. Therefore, for each pair of formulas, we first computed all possible transformations of each formula to obtain a set of all its variants. Then, we compared the formula structures in the two sets to determine if the two formulas shared a structure. If so, we checked whether their symbols, with subscripts removed, were identical. If so, we considered these two formulas to be mergeable. When merging, we retained the parameter with the shorter name of the two. After merging based on such symbolic rules, we reduced the number of formulas in the formula database from 12,906 to 1,163.
Semantics based merging
In the previous step, the semantic information in the parameter names was not used. This led us to perform merges grounded on the semantics of parameter names. For example, two formulas that were not merged during the symbolic merging stage, “[density] = [mass] / [volume]” and “[density of water] = [mass of water] / [volume of water]”, should actually be merged. We identified such mergeable formulas based on the semantic information in the parameter names, e.g., “density” and “density of water” are semantically similar. Specifically, for formulas with identical structures, we tokenized [^3] We used jieba: https://github.com/fxsjy/jieba. each pair of their corresponding parameter names into two sets of words. When the two sets overlapped, the two parameters were considered to have a semantic connection and the two formulas became candidates for merging. Using this approach, we identified a set of pairs of potentially mergeable formulas and then consulted the LLM for a detailed evaluation of each pair. The prompt is provided in Appendix A.1.3. After this step, the number of formulas in the formula database was reduced from 1,163 to 439.
Manual review and error correction
Upon completing the aforementioned merging process, we manually inspected the accuracy of the results, rectified the instances where errors occurred during merging, and manually merged formulas that were overlooked by the LLM. In this process, two human volunteers cross-validated the results of manual review and correction. Finally, we obtained a consolidated formula database consisting of 272 distinct formulas.
3.4 Question Alteration
The original questions in FormulaReasoning were collected from publicly accessible online sources. Considering that LLMs were typically trained on massive-scale corpora that potentially encompassed these publicly available resources, there existed a non-negligible risk of data contamination and label leakage. To mitigate this potential bias and construct a more rigorous benchmark, we implemented systematic strategies for question alteration. Specifically, we employed inverse transformation (You et al., 2024) and parameter noise injection (Wei and Zou, 2019) to generate a modified version for each original question while preserving its fundamental reasoning nature.
Inverse transformation
[…]
Parameter noise injection
We randomly selected a parameter from the entire set of parameters in all questions and appended both its parameter name and its numerical value to the question as a distractor. To ensure that the added parameter would not affect the solvability of the question (e.g., not introducing two conflicting values for the same physical quantity), we ensured that the added parameter was different from all the parameters involved in answering the original question.
To verify that the modified questions were solvable, we used SymPy [^4] https://www.sympy.org/ to solve the equations. During the verification process, some equations could not be solved (e.g., SymPy cannot handle “%” and “1000kg/t”), so we removed these questions. The final number of the remaining questions is 5,324.
3.5 English Version of FormulaReasoning
We employed the LLM to translate Chinese questions, explanations, and formula database into English. The translation prompts are provided in Appendix A.1.4. To evaluate translation quality, we manually translated a sample of 50 questions and calculated the perplexity of the LLM-translated version and the manually translated version using Qwen models for comparison. We computed the perplexity of LLM and human translations using the Qwen2.5-14B-Instruct model, resulting in scores of $6.28$ and $7.26$, respectively. These findings demonstrate that the current English version maintains a reasonable level of fluency.
We also manually compared the Chinese and translated English versions of the 50 sampled questions. We found that no question had translation errors that would affect question answering.
We provide both Chinese and English versions of FormulaReasoning.
4 Experimental Setup
We implemented a wide range of reasoning approaches for handling the questions within FormulaReasoning, with a focus on understanding how different methodological choices impact performance on formula-based numerical reasoning tasks. Our investigation includes analyzing various prompting strategies for language models, developing specialized fine-tuning techniques, and incorporating external knowledge through retrieval-augmented methods. Specifically, we developed two complementary approaches to enhance reasoning capabilities: one that decomposed the reasoning process into formula generation, parameter extraction, and numerical calculation, and another that leveraged data augmentation techniques. We also explored preference learning to refine reasoning processes. This comprehensive experimental setup allows us to compare the strengths and limitations of different architectural and methodological choices for formula-based reasoning.
4.1 Dataset Split
We divided FormulaReasoning into three subsets: training, HoF (Homologous Formulas) test, and HeF (Heterologous Formulas) test, comprising 4,524, 413, and 387 questions, respectively. All formulas in the HoF test set appeared in the training set, while in the HeF test set, each question required at least one formula not seen in the training set. This division was to evaluate the generalizability of fine-tuned models on new formulas.
4.2 Evaluated Methods
LLMs
Following Kojima et al. (2022), we incorporated “Let’s think step by step” (i.e. CoT) into a zero-shot prompt to guide LLMs in generating reasoning steps. The prompt is in Appendix A.2.1.
We conducted experiments across a diverse spectrum of models including cost friendly small-scale models including GLM-4.5-Air (GLM et al., 2024), ERNIE-4.5-21B-A3B Baidu-ERNIE-Team (2025), Gemma-3N-E4B Gemma (2025), Llama-3.3-8B-Instruct (Meta, 2024), Mistral-Small-3.2-24b-Instruct (Mistral, 2025), GPT-oss-20B (Agarwal et al., 2025), Gemini-2.5-Flash-Lite (Comanici et al., 2025) and Seed-OSS-36B-Instruct Team (2025), alongside high-cost LLMs such as GPT-4o (OpenAI, 2024), GLM-4-plus (GLM et al., 2024), Qwen3-235B-A22B-Thinking-2507 (Yang et al., 2025), DeepSeek-R1 (DeepSeek-AI et al., 2025), o3-mini (OpenAI, 2025b), and GPT-5 (OpenAI, 2025a). This comprehensive evaluation allowed us to assess capabilities across different model scales while identifying persistent challenges in formula reasoning.
We also compared CoT with Program of Thought (PoT) Chen et al. (2023). In PoT, we used a Python interpreter to execute the code and obtain an answer.
Formula Retriever
We trained a formula retriever on the training set. Specifically, we encoded each question using the Chinese-BERT-wwm-base model (Devlin et al., 2019; Cui et al., 2021) to obtain the CLS vector of the question. The formulas in the formula database were not encoded in this way, as they contained structured information so that two textually similar formulas could have entirely different meanings (e.g., $z_{1}=x\times y$ and $z_{2}=x/y$ ). Therefore, each formula was instead represented by a randomly initialized vector that was updated during training. We calculated the cosine score between the question vector and the formula vector. The retriever was then trained with in-batch negatives and contrastive learning loss Gao et al. (2021). During inference, for each question in the HoF test, we retrieved the top five formulas with the highest scores and included them in the prompt to augment LLM generation. More details are provided in Appendix A.2.2.
Supervised Fine-Tuned Models
We found that directly prompting models having fewer than 8B parameters did not produce satisfactory results, so we performed supervised fine-tuning of small models. We conducted experiments on Qwen2.5-Math-7B-Instruct and Qwen2.5-Math-1.5B-Instruct.
Unlike large models, small models struggled with numerical extraction and calculation. To enhance their reasoning capabilities, we developed:
  • CoT-Supervised Fine-Tuning (CoT-SFT): A three-step process where the model generates relevant formulas, extracts parameter values and units from questions, and uses Numbat for final calculation.
  • Data Augmentation (DA): Using Qwen-max to generate and verify new training examples with correct reasoning steps. Details are provided in Appendix A.2.3.
Direct Preference Optimization (DPO)
We used the preference data derived in Appendix A.2.4 to experiment on DPO. In the generation process, we used DeepSeek-R1-Distill-Qwen-7B as the generation model, and used Qwen2.5-Math-PRM-7B as the process reward model (PRM).
The implementation details are provided in Appendix A.2.5.
4.3 Evaluation Metrics
We utilized Numbat to evaluate the prediction generated by each model against the gold-standard answer. A prediction was deemed correct if the relative error, (prediction - gold) / gold, was less than 1%. We used accuracy, which is the proportion of questions correctly answered, as our metric.
We also used PRM score to evaluate the reasoning steps generated by LLMs. There were two evaluation settings: one-step and multi-step. One-step evaluation scored the overall output using PRM, whereas multi-step evaluation divided the output into multiple steps, scored each individually, and took the average as the final score. Following the setting of DPO mentioned above, here we again used Qwen2.5-Math-PRM-7B as the PRM.
5 Experimental Results
We present the evaluation results on the Chinese version of FormulaReasoning. The results on the English version are in Appendix A.3.1, which are consistent with those on the Chinese version. The results of PoT are in A.3.2, not better than CoT.
Table 4: Accuracy of LLMs with CoT prompts.
Model HoF test HeF test Avg. GLM-4.5-Air 59.32 61.50 60.37 ERNIE-4.5-21B-A3B 66.34 67.96 67.12 Gemma-3n-E4B-it 14.04 14.47 14.25 Llama-3.3-8b-Instruct 8.47 9.56 9.00 Mistral-Small-3.2-24b-Instruct 74.33 72.35 73.37 GPT-oss-20B 85.96 82.95 84.50 Gemini-2.5-Flash-Lite 81.11 81.40 81.25 Seed-OSS-36B-Instruct 81.84 84.24 83.00 GPT-4o 69.49 65.37 67.50 GLM-4-plus 61.74 63.57 62.63 Qwen3-235B-A22B-Thinking-2507 80.63 83.98 82.25 DeepSeek-R1 88.62 91.47 90.00 o3-mini 92.01 89.66 90.87 GPT-5 92.01 92.76 92.37 Human 93.49 90.47 92.03
5.1 Results of LLMs with CoT Prompts
We evaluated a diverse range of LLMs across different parameter scales on FormulaReasoning. Our analysis reveals that model performance correlates with both parameter scale and architectural design. As shown in Table 4, large-scale models (e.g., DeepSeek-R1 and o3-mini) achieved competitive results with accuracies exceeding 88% on both HoF and HeF metrics. However, substantial performance gaps existed for smaller-scale models, where architectures like Llama-3.3-8B and Gemma-3N-E4B-it exhibit accuracies below 15%, while even mid-sized models like GLM-4-Air only reached around 60%. Notably, despite their impressive capabilities, even advanced large-scale models such as GPT-4o (67.50%), GLM-4-plus (62.63%), and Qwen3-235B (82.25%) still trailed behind human performance (92.03%). This significant variance suggests that while parameter scaling yields notable improvements, there remains considerable optimization potential for resource-efficient models. Future work should investigate architectural adaptations and training strategies specifically tailored for compact models, which are critical for practical deployment scenarios where computational resources are constrained. The observed performance spectrum underscores the importance of scale-aware evaluation in formula-based reasoning research.
We provide an error analysis in Appendix A.4.
PRM scores to assess the reasoning steps generated by LLMs are included in Appendix A.3.3.
5.2 Results of LLMs with Formula Retriever
We found that incorporating retrieved formulas into the prompts improved the performance of LLMs. For GLM-4-Air, the accuracy increased by 1.94 from 65.13 to 67.07 on the HoF. Note that it would be meaningless to carry out this experiment on the HeF test set where the required formulas were not included when training the retriever.
Our investigation found that the top five formulas retrieved often included irrelevant ones, as the number of formulas required varied for different questions. The presence of extraneous formulas might introduce noise, suggesting considerable room for improvement in structured formula retrieval.
5.3 Results of Supervised Fine-Tuned Models
The results of supervised fine-tuned small models are included in Appendix A.3.4. Both CoT-SFT and DA helped improve accuracy. In particular, Qwen2.5-Math-7B-Instruct with CoT-SFT outperformed GPT-4o, demonstrating the effectiveness of our fine-tuning method: delegating numerical calculations to a calculator and focusing on CoT reasoning.
5.4 Results of DPO (including Results on Other Numerical Reasoning Datasets)
As shown in Appendix A.3.5, DPO with the preference data derived from the training set of FormulaReasoning generally improved the accuracy of 1.5B–7B models on both FormulaReasoning and other numerical reasoning datasets: GSM8K (Cobbe et al., 2021a), MATH Hendrycks et al. (2021), SVAMP Patel et al. (2021), and GaoKao2023-en Chen et al. (2025). These models exhibited much lower accuracy on FormulaReasoning than those datasets. The results underscore the distinct characteristics of our dataset compared to other numerical reasoning benchmarks and its additional value as a potential DPO resource.
6 Conclusion
We introduced FormulaReasoning, a bilingual dataset for formula-based numerical reasoning. For each question, we annotated its complex reasoning steps with normalized physics formulas. We constructed a consolidated formula database after merging equivalent formulas, serving as an external knowledge base to be used in RAG. We used our dataset to evaluate LLMs of various sizes, RAG, fine-tuned small models, and DPO methods, revealing significant performance variations across model scales. Our findings highlight that while larger models show promising performance, smaller-scale LLMs still face significant challenges in formula-based numerical reasoning, indicating opportunities for further advancements in multi-step reasoning guided by domain knowledge.
In future work, we plan to use the formular knowledge in FormulaReasoning to improve the numerical reasoning capabilities of LLMs. Possible approaches include enhancing reasoning abilities through knowledge-driven methods or reinforcement learning methods based on formula feedback.
Limitations
One limitation of our work is that although FormulaReasoning provides step-level supervision information (i.e., formulas), due to the diversity of reasoning paths and formula expressions, we only used PRM to perform step-level process evaluation. Another limitation is that our dataset is focused on physics. We chose junior high school physics because it is not too hard to be understood by ordinary people, which benefited our annotation and evaluation efforts. It is possible to explore formula-based numerical reasoning in other domains such as chemistry and engineering.
Appendix A Appendix
A.1 Appendix for Dataset Construction
A.1.1 Prompts for Formula Normalization
The formula normalization process followed a coarse-to-fine annotation process as described in Section 3.2. In the coarse-grained annotation phase, the natural language explanations were normalized and the associated parameters were extracted from these explanations. These prompts are provided in Figures 2 and 3. The fine-grained annotation phase focused on error correction, which was divided into three specific categories: input errors, where the parameters mentioned in the explanation were absent from the question; calculation errors, which occurred when the Numbat calculator reported an error during the computation process; and output errors, where the final calculated answer was incorrect. We provide a prompt for correcting the calculation errors as an example in Figure 4; the prompts for the other two error types are provided in our GitHub repository. The entire normalization procedure used a few-shot prompt with six examples.
Figure 2: Prompt for explanation normalization.
Figure 3: Prompt for parameter extraction.
Figure 4: Prompt for correcting calculation errors.
A.1.2 Examples of Removed Questions
The questions that remained incorrect despite multiple attempts by the LLM were of notably poor quality, e.g. missing important reasoning steps or having an incorrect reference answer. We provide an example of these questions in Figure 5.
Figure 5: An example of removed question.
A.1.3 Prompt for Semantics Based Merging
Semantics based merging primarily employed the LLM to comprehend formulas, determined if two formulas were semantically equivalent, and subsequently determined whether they could be merged into a single formula. The prompt for this procedure is illustrated in Figure 6. This approach ensured that nuanced meanings embedded within formulas were accurately captured and evaluated for potential merging, thereby enhancing the quality of the formula database.
Figure 6: Prompt for semantics based merging.
A.1.4 Prompt for Dataset Translation
The prompts for translating the dataset are provided in Figure 7.
Figure 7: Prompts for translating questions (top) and formulas (bottom).
A.2 Appendix for Experimental Setup
A.2.1 Prompt for Evaluated LLMs
The prompt is provided in Figure 8.
Figure 8: Prompt for evaluated LLMs.
A.2.2 Formula Retriever
Implementation
Let the number of formulas in the formula database be $N$. During training, we randomly initialized a matrix $\mathbf{F}\in\mathbb{R}^{N\times d}$, where $d$ is the hidden size and the $i$-th row in $\mathbf{F}$ represented the initial representation of the $i$-th formula in formula database. We denoted a batch of questions with a batch size of $B$ as $Q=\{q_{1},q_{2},...,q_{B}\}$. The indices of the gold-standard formulas corresponding to these $B$ questions were denoted as $L=\{l_{1},l_{2},\cdots,l_{B}\}$ (i.e. the label of $q_{i}$ is $l_{i}$, where $1\leq i\leq B$).
BERT was utilized to encode each question,
$\mathbf{h}^{i}_{cls},\mathbf{h}^{i}_{1},\cdots=\mathtt{BERT}(q_{i}),1\leq i\leq B.$ (1)
Subsequently, we took the CLS vector $\mathbf{h}^{i}_{cls}$ as the representation for the $i$-th question.
We utilized in-batch negatives and contrastive learning loss,
$\mathcal{L}=-\frac{1}{B}\sum_{1\leq i\leq B}\log\frac{\exp(\cos(\mathbf{h}^{i}_{cls},\mathbf{F}_{l_{i}}))}{\sum_{1\leq j\leq B}\exp(\cos(\mathbf{h}^{i}_{cls},\mathbf{F}_{l_{j}}))}.$ (2)
Each question might correspond to multiple correct formulas, and we ensured that the same question did not appear twice in the same batch when loading the data. Based on the implementation of Chinese-BERT-wwm-base, we tested the retrieval performance on the HoF test set and found that Recall@5 reached 82.46%.
Prompt for Evaluated LLMs with Formula Retriever
For each question, we included the top-5 retrieved formulas in the prompt provided in Figure 9.
Figure 9: Prompt for evaluated LLMs with formula retriever.
A.2.3 Data Augmentation for Supervised Fine-Tuned Models
Several studies (Ding et al., 2024; Zheng et al., 2023; Whitehouse et al., 2023) use LLM for data augmentation, which mainly focus on daily conversations or sentiment analysis and do not require rigorous numerical calculations. Some research (Shum et al., 2023) on data augmentation that involves numerical calculations employs LLM to generate solutions for existing questions to aid training, rather than to generate new questions. In contrast to these approaches, we generated complete questions that involve numerical calculations (particularly focusing on formulas), along with automatic improvement and selection to ensure data quality.
We divided the data generation process into the following steps. First, we randomly generated 17,000 prompts. Each prompt was obtained by stacking five question-explanation pairs sampled from the training set. At the end of the prompt, the LLM was required to generate the sixth question-explanation pair. Second, we normalized the formulas generated. Except for the absence of manual review, the remaining steps were consistent with those in Section 3.2. Finally, we utilized Numbat to check whether the calculation process in the data generated by the LLM was correct, and discarded the data with incorrect calculations. After the above steps, we finally retained about 2,500 questions.
We found that mixing the newly generated data into the original training set did not always bring positive improvement, perhaps because the newly generated data did not undergo manual review. We found that randomly selecting a small portion of the newly generated data could improve the performance of the model. We set the mixing ratio to 10% which is an experiential proportion. We fine-tuned each model using the augmented dataset. After training for a fixed number of epochs (50 epochs), we selected the final checkpoints as DA models.
A.2.4 Preference Data Generation
We applied the Monte Carlo Tree Search (MCTS) method to the unaltered questions in the training set of FormulaReasoning to generate preference data for DPO. It used the reasoning steps as nodes, and used the generation model to expand nodes. Each node had one more reasoning step than its parent node. The root node had no reasoning step, only containing the original question. The termination state was reached when the termination token was generated or the current question was solved. We used the PRM to guide the generation. The PRM score was used as a reward to update the nodes.
After generation, a tree was created for each sample. Each node contained the reasoning steps generated up to that node and the correctness score of the generated steps. We identified the best and worst paths in the tree. Each path was scored by the average score from the root node to the terminated leaf node. Limited by the number of iterations, some leaf nodes might not represent termination and their corresponding paths were not used.
To consider the inclusion of formula knowledge, we further screened the generated data. For each sample, we examined its best and worst paths. If there was no significant formula knowledge on these paths, we would discard this sample. We used a rule-based pattern matching method to identify formula knowledge. This pattern was consistent with the example given in the generation prompt used in MCTS, which was an expression consisting of "[{parameter}]".
A.2.5 Implementation Details
We accessed DeepSeek-R1 through the DeepSeek API [^5] https://api-docs.deepseek.com/, Qwen series models through the Alibaba Bailian platform [^6] https://bailian.console.aliyun.com/, and other models through OpenRouter [^7] https://openrouter.ai/ with their default hyper-parameters. Qwen2.5-Math-7B/1.5B-Instruct were fine-tuned on NVIDIA GeForce RTX 3090 GPUs with Huggingface Transformers and Pytorch 2.3, where we used a learning rate of 1e-4 and a batch size of 8 to fine-tune them. For Qwen2.5-Math-7B-Instruct model, we fine-tune it with LoRA (Hu et al., 2021) with r=8, alpha=32. For Qwen2.5-Math-1.5B-Instruct model, we fine-tune it with full fine-tuning. We tested the models after training for 3 epochs which is enough for training. We set temperature=0 for direct answer generation and set the maximum output length to 1,024. In the case of CoT-SFT, which directly outputted formulas along with numerical values and units of parameters, we set temperature=0 to obtain the response with the highest probability and set the maximum output length to 512 to obtain formulas and parameters. Training Qwen2.5-Math-7B-Instruct(lora) and Qwen2.5-Math-1.5B-Instruct(full) in CoT-SFT used 3 and 8 hours, respectively.
A.3 Appendix for Experimental Results
A.3.1 English Version of FormulaReasoning
The evaluation results for LLMs with CoT prompts on the English FormulaReasoning dataset are presented in Table 5. Large-scale models (e.g., GPT-5 and DeepSeek-R1) achieved competitive results with accuracies exceeding 87% on both HoF and HeF metrics. However, substantial performance gaps existed for smaller-scale models, where architectures like Gemma-3n-E4B-it and Llama-3.3-8b-Instruct exhibited accuracies below 35%, while even mid-sized models like GLM-4.5-Air only reached around 73%. We found that Gemma-3n and Qwen3-235B showed performance drops compared to their results on the Chinese version due to instruction-following issues in the English version, where these models failed to output required units as specified in the prompt.
Notably, despite their impressive capabilities, even the best model GPT-5 (88.63%) still trailed behind human performance (92.03% as reported in the main text), falling short by 3.4 percentage points. Other advanced models such as GPT-4o (75.38%) and GLM-4-plus (72.37%) showed even larger gaps with human performance.
This significant variance suggests that while parameter scaling yields notable improvements, there remains considerable optimization potential for resource-efficient models. Future work should investigate architectural adaptations and training strategies specifically tailored for compact models, which are critical for practical deployment scenarios where computational resources are constrained. The observed performance spectrum underscores the importance of scale-aware evaluation in formula-based reasoning research.
Table 5: Accuracy of LLMs with CoT prompts on English FormulaReasoning.
Model HoF test HeF test Avg. GLM-4.5-Air 72.64 74.42 73.53 ERNIE-4.5-21B-A3B 63.20 56.85 60.13 Gemma-3n-E4B-it 2.66 2.84 2.75 Llama-3.3-8b-Instruct 32.45 31.52 32.00 Mistral-Small-3.2-24b-Instruct 70.70 68.99 69.87 GPT-oss-20B 86.20 84.75 85.48 Gemini-2.5-Flash-Lite 75.30 76.49 75.88 Seed-OSS-36B-Instruct 75.30 71.58 73.50 GPT-4o 75.79 74.94 75.38 GLM-4-plus 75.79 68.73 72.37 Qwen3-235B-A22B-Thinking-2507 62.47 66.67 64.50 DeepSeek-R1 87.17 87.60 87.38 o3-mini 84.02 82.17 83.13 GPT-5 88.38 88.89 88.63
A.3.2 PoT Prompts
The results of GPT-4o using CoT compared to PoT are shown in Table 6. CoT consistently outperformed PoT. One possible reason was that the effectiveness of PoT could be influenced by a hybrid of the language and coding capabilities of LLMs, but that of CoT was only related to the language capabilities. Moreover, for LLMs with strong calculation capabilities, the advantage of using Python to perform numerical calculations might be eroded. Therefore, on FormulaReasoning, the bottleneck for current LLMs was mainly in their ability to apply formulas, rather than in calculation.
Table 6: Accuracy of LLMs with CoT and PoT prompts.
Model HoF test HeF test Avg.
GPT-4o (CoT) 69.49 65.37 67.50
GPT-4o (PoT) 54.72 41.34 48.25
A.3.3 PRM Scores
Tables 7 and 8 show the PRM scores in one-step and multi-step settings, respectively.
Table 7: PRM score (one-step) of LLMs with CoT prompts.
Model HoF test HeF test Avg. GLM-4.5-Air 0.9301 0.9436 0.9369 ERNIE-4.5-21B-A3B 0.7753 0.7582 0.7668 Gemma-3n-E4B-it 0.3695 0.3660 0.3678 Llama-3.3-8b-Instruct 0.2984 0.2408 0.2696 Mistral-Small-3.2-24b-Instruct 0.8576 0.8133 0.8355 GPT-oss-20B 0.5704 0.5203 0.5454 Gemini-2.5-Flash-Lite 0.5821 0.5818 0.5820 Seed-OSS-36B-Instruct 0.8766 0.8764 0.8765 GPT-4o 0.6150 0.5583 0.5867 GLM-4-plus 0.6728 0.6591 0.6660 Qwen3-235B-A22B-Thinking-2507 0.6502 0.6476 0.6489 DeepSeek-R1 0.7395 0.7393 0.7394 o3-mini 0.7537 0.7556 0.7547 GPT-5 0.8196 0.8249 0.8223
Table 8: PRM score (multi-step) of LLMs with CoT prompts.
Model HoF test HeF test Avg. GLM-4.5-Air 0.9336 0.9438 0.9387 ERNIE-4.5-21B-A3B 0.8695 0.8711 0.8703 Gemma-3n-E4B-it 0.7366 0.7360 0.7363 Llama-3.3-8b-Instruct 0.5507 0.5288 0.5398 Mistral-Small-3.2-24b-Instruct 0.9427 0.9438 0.9433 GPT-oss-20B 0.5738 0.5266 0.5502 Gemini-2.5-Flash-Lite 0.8621 0.8719 0.8670 Seed-OSS-36B-Instruct 0.8789 0.8779 0.8784 GPT-4o 0.8655 0.8444 0.8550 GLM-4-plus 0.8726 0.8738 0.8732 Qwen3-235B-A22B-Thinking-2507 0.7946 0.7996 0.7971 DeepSeek-R1 0.7603 0.7535 0.7569 o3-mini 0.8104 0.8032 0.8068 GPT-5 0.8196 0.8248 0.8222
Table 9: Accuracy of supervised fine-tuned models.
Model HoF test HeF test Avg.
Qwen2.5-Math-7B-Instruct 65.62 58.91 62.27
   + CoT-SFT 72.64 +7.02+7.02 65.63 +6.72+6.72 69.14 +6.87+6.87
   + DA 67.55 +1.93+1.93 62.53 +3.62+3.62 65.04 +2.77+2.77
Qwen2.5-Math-1.5B-Instruct 64.41 55.56 59.99
   + CoT-SFT 66.34 +1.93+1.93 53.75 1.81-1.81 60.05 +0.06+0.06
   + DA 66.59 +2.18+2.18 59.17 +3.61+3.61 62.88 +2.89+2.89
A.3.4 Supervised Fine-Tuned Models
Table 9 shows the results for the supervised fine-tuned models, with and without CoT-SFT or DA. In most settings, both models achieved higher scores on the HoF test set than on HeF, yet they still exhibited considerable performance on the latter. This indicated that the unseen formulas influenced the performance of the models, though they still demonstrated a level of generalizability.
Interestingly, by adding CoT-SFT, Qwen2.5-Math-7B-Instruct with a mere parameter count of 7B achieved a performance better than GPT-4o, although such a comparison might not be entirely fair. The results revealed that focusing on reasoning with CoT while delegating numerical calculations to a calculator greatly improved the performance of small models, given their limited calculation capacity. Data augmentation also lightly improved the reasoning capacity of small models.
Table 10: Accuracy of DPO on FormulaReasoning and other datasets.
Model GSM8K MATH SVAMP Gaokao2023-en FormulaReasoning
DeepSeek-R1-Distill-Qwen-7B 87.7 58.8 87.7 55.3 34.0
   +DPO 85.8 1.9-1.9 59.7 +0.9+0.9 88.1 +0.4+0.4 55.8 +0.5+0.5 37.0 +3.0+3.0
DeepSeek-R1-Distill-Qwen-1.5B 79.8 68.2 85.7 55.3 7.23
   +DPO 81.3 +1.5+1.5 68.0 0.2-0.2 86.3 +0.6+0.6 55.1 0.2-0.2 10.47 +3.24+3.24
Qwen2.5-Math-1.5B-Instruct 85.6 76.1 93.3 66.0 4.94
   +DPO 85.8 +0.2+0.2 76.5 +0.4+0.4 93.7 +0.4+0.4 65.2 0.8-0.8 7.52 +2.58+2.58
Qwen2.5-1.5B-Instruct 72.9 55.0 85.5 47.5 5.94
   +DPO 73.1 +0.2+0.2 56.2 +1.2+1.2 86.6 +1.1+1.1 47.8 +0.3+0.3 6.57 +0.63+0.63
A.3.5 DPO
We employed the preference data derived from the training set of FormulaReasoning to perform DPO with several small models on not only the test sets of FormulaReasoning but also other numerical reasoning datasets. Table 10 shows the results. DPO improved accuracy on FormulaReasoning and also improved accuracy in most cases on other datasets, showing an additional value of the preference data derived from our dataset. Notably, although these models exhibited good performance on other datasets, their low accuracy on FormulaReasoning indicated that they did not truly master formula-based reasoning required in the physics domain, characterizing FormulaReasoning as a valuable dataset presenting unique challenges for LLMs.
A.4 Error Analysis
FormulaReasoning posed challenges to existing models in terms of formula selection and numerical calculation, including unit calculation and arithmetic calculation. As illustrated in Figure 10(a), the incorrect use of formulas was the main type of error. Another type of error was parameter noise, as illustrated in Figure 10(b). In terms of the performance of the models, we found that the models that achieved higher accuracy were less affected by the parameter noise injected into the questions (Section 3.4). For example, among the questions to which DeepSeek-R1 did not provide a correct answer, approximately 14% were influenced by the parameter noise, while for Qwen-max, the number rose to 28%. This indicated that our question alteration imposed further requirements on the model’s ability to correctly apply formulas and perform accurate reasoning.
(a) An error case caused by formula error.
(b) An error case caused by parameter noise.
Figure 10: Examples of errors.